LHC 2010 - latest news

Week 10: Machine Coordinators: Ralph Assmann - Mike Lamont

Latest planning outline (10-03-10)

Preparing ramp.
Then pre-cycle. Then injection & beam dump studies.

Technical stop next Monday andTuesday!

This week (so far) - Availability 110h40. Unplanned maintenance: 28h20. Planned maintenance: 29h00.

Day Short summary Link to report from 8:30 and 17:00 meetings
Sunday 14th March
  • 07:00: Pre-cycle in preparation of ramp.
  • 09:00: Preparing ramp to 1.18 TeV with beam.
  • 12:30: Ramp to 1.18 TeV. Measurements at 1.18 TeV.
  • 15:00: Dump 1.18 TeV fill. Cycle.
  • 19:00: Injection and beam dump studies.
  • Availability: 24h00. Unplanned maintenance: 0h00. Planned maintenance: 0h00.
Slides from 9:00 meeting
Saturday 13th March
  • 07:00: Studies with aperture kicker started. Kicking demonstrated.
  • 11:00: Beam setup, loss optimization, 2 RF cavities restarted. LHCb magnets turned on. Orbit corrected.
  • 16:00: Beam tuning of pilot bunch: lifetime > 125h both beams. Emittance.
  • 17:00: Injection of several single bunch intensities (pilot, fat pilot, half nominal): 5e9-6e10.
  • 21:00: LHC beam quality checks in SPS and TL.
  • 22:30: Reinjection LHC. Halo stability check.
  • 23:30: Aperture studies.
  • Availability: 24h00. Unplanned maintenance: 0h00. Planned maintenance: 0h00.
Slides from 9:00 meeting
Friday 12th March
  • 02:00: Machine closed.
  • 09:00: Cryogenics all OK. Preparing pre-cycle.
  • 10:00: Pre-cycle started.
  • 11:30: Pre-cycle finished.
  • 14:00: Beam back. Orbit corrections, injection check, tune adjustments.
  • 15:00: PIC on 60A power converters.
  • 16:00: Beam back. Chromaticity adjustments. Beta beat. Orbit feedback tests.
  • 21:00: Start of ramp with beam to 1.18 TeV. Arrive very nicely at top at 21:11.
  • 23:00: Aperture measurements.
  • Availability: 14h00. Unplanned maintenance: 1h00. Planned maintenance: 9h00.
Slides from 8:30 meeting
Thursday 11th March
  • 00:00: Beam commissioning of transverse damper.
  • 04:00: Start of cryo work.
  • 07:00: Planned access to the machine (nQPS etc).
  • Availability: 04h00. Unplanned maintenance: 0h00. Planned maintenance: 20h00.
Slides from 8:30 meeting
Wednesday 10th March
  • 00:20: Second ramp without beam to 1.18 TeV. BETS analyzing. Then ramp-down precycle combo.
  • 03:00: Beam back after ramp tests without beam. Tunes, beta beat, chromaticity measurements and adjustments. b3 knob calibration.
  • 08:00: Beam commissioning work on the RF system and tune feedback.
  • 12:00: MPS tests.
  • 16:00: Injection studies.
  • Availability: 24h00. Unplanned maintenance: 0h00. Planned maintenance: 0h00.

Slides from 8:30 meeting

Slides from LMC report

Tuesday 9th March
  • Orbit feedback tests curtailed by cryogenics controls problem at 05:00
  • Seamless merged with current lead problem in S45 at around 07:00
  • 16:00: Both beams are back. Checking and correcting tunes & orbit.
  • 17:00: Collimation tests starting. Loss maps for efficiency & passive protection. Grazing events to CMS.
  • 20:00: Prepare for ramp test without beam. 21:42 machine at 1.18 TeV without beam.
  • 22:00: Ramp-down precycle combo.
  • Availability: 13h00. Unplanned maintenance: 11h00. Planned maintenance: 0h00.
Slides from 8:30 meeting
Mon 8th March 2010
  • Since 01h40: Recovery from emergency stop due to switch failure.
  • 14h00: Access finished, but beam imminent warning problems.
  • 16h45: All sectors precycled or precycling: quench of Q7L5 and some minor issues.
  • 18h00: machine ready for beam, preparing injection, resolving interlock and BPM issues.
  • 20h30: both beams back in LHC, happily circulating. Start injection and dump protection setting-up. Then over-night test Q and orbit feedback.
  • Availability: 7h40. Unplanned maintenance: 16h20. Planned maintenance: 0h00.
Presentation at the 8:30 meeting